OpenStudy (anonymous):

Evaluate the expression for the given value of the variable(s) |-4b-8| + |-1-b^2| + 2b^3 ; b=-2

OpenStudy (anonymous):

l-4(-2)-8l + l-1-2^2l + 2(-2)^3= 16+3+10=29